OpenStudy (veeveeniko):

How would I write the equation of a circle with a center at (0, 0) and a radius of 5?

OpenStudy (anonymous):

this is the fomrula for it (x-a)2 + (y-b)2 = r2

OpenStudy (anonymous):

(h-0)^2 + (k-0)^2 = 25

OpenStudy (veeveeniko):

I don't really understand still.

OpenStudy (anonymous):

or x^2 + y^2 = 25

OpenStudy (anonymous):

As an example, let us put some values to a, b and r and then expand it Start with: (x-a)2 + (y-b)2 = r2 Set (for example) a=1, b=2, c=3: (x-1)2 + (y-2)2 = 32 Expand: x2 - 2x + 1 + y2 - 4y + 4 = 9 Gather like terms: x2 + y2 - 2x - 4y + 1 + 4 - 9 = 0 And we end up with this: x2 + y2 - 2x - 4y - 4 = 0

OpenStudy (veeveeniko):

What I meant was how were you able to determine what ab, b and c is from the information i provided?

OpenStudy (veeveeniko):

a, b and c*

OpenStudy (anonymous):

a is x1 b is y1

OpenStudy (anonymous):

c is radious squared

OpenStudy (veeveeniko):

oh okay, so for my problem x1=0 and y1=0? and c would be 25

OpenStudy (anonymous):

Ok so the circle formula is (x-h)^2 + (y-k)^2 = r^2 The midpoint will be (h, k) with radius (r).

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

listen to cosm also hes correct

OpenStudy (veeveeniko):

alright that is very helpful thank you both

OpenStudy (anonymous):

heres a link also for further halp with this http://www.mathsisfun.com/algebra/circle-equations.html

OpenStudy (anonymous):

x^2+y^2=25